when leonard is five years oldertan double his present age he will be three times as old as he was a year ago. how old is he?
Answer by ankor@dixie-net.com(22740)

when Leonard is five years older than double his present age he will be
three times as old as he was a year ago. how old is he?
;
Let a = his present age
:
2a + 5 = 3(a - 1)
:
2a + 5 = 3a - 3
:
2a - 3a = -3 - 5
:
-a = -8
:
a = 8 yrs old is his present age
:
:
Check solution in original equation
2(8) + 5 = 3(7)